[Brahms-l] Shift report 20040125 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sun Jan 25 2004 - 07:32:49 EST
------------------------------------------------------------
Date: Sun Jan 25 07:32:49 2004, Shift: 00:00-08:00
Supervisor: Ionut ARSENE, with: 
Log book#15 pages 42 - 42

Shift Summary:
------------------------------------------------------------
1. Beam for almost the entire shift. No problems.

2. TPC and non-TPC pedestals (9849 and 9848, respectively)
   The setting D5@B2255 with MRS at 40deg is done. (run 9847)
   I changed the polarity in D5 to A, same current.
   The next runs are for this setting: 9850, 9851, 9852, 9853. 
FS settings for the entire shift: 10deg with D1@B1219. This setting is DONE too.

3. No problems with the detectors.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 25 07:33:02 2004

This archive was generated by hypermail 2.1.8 : Sun Jan 25 2004 - 07:33:21 EST